If A= 4m and B= 3m for two vectors and the the angle between them is 450, Calculate AxB
1
Expert's answer
2021-02-19T18:58:10-0500

The vector product of two vectors can be found as follows:


"\vec{A}\times\vec{B}=|\vec{A}|\cdot|\vec{B}|sin\theta,""\vec{A}\times\vec{B}=4\ m\cdot3\ m\cdot sin45^{\circ}=8.48\ m^2."